i am currentley revising for my gcse and i am stuck on algebra. the school have given me a a word dotted paper and told me to work out the area by drawing the shapes.There is a word on the paper i do not understand the word is trapezium. what does that word mean can you help me please thank you.
Answer by venugopalramana(3286)

i am currentley revising for my gcse and i am stuck on algebra. the school have given me a a word dotted paper and told me to work out the area by drawing the shapes.There is a word on the paper i do not understand the word is trapezium. what does that word mean can you help me please thank you.
SEE THE FOLLOWING EXAMPLES TO KNOW THE TRAPEZIUM..IT IS A QUADRILATERAL WITH 2 OPPOSITE SIDES PARALLEL
IF YOU WANT MORE OR IN DOUBT COME BACK.
see the following example and try
SEE THE TRAPEZOID ....ABCD below
AREA OF TRAPEZOID IS GIVEN BY
AREA=(1/2)*ALTITUDE*(SUM OF PARALLEL SIDES)=(1/2)*DP*(AB+CD)
WHERE ALTITUDE IS THE PERPENDICULAR DISTANCE BETWEEN PARALLEL SIDES =DP
PARALLEL SIDES ARE AB AND CD.
WE HAVE AB=40
BC=AD=20
ANGLE DAP=63
DAP IS A RIGHT ANGLED TRIANGLE...DP/DA=SIN(63)
DP=DA*SIN(63)=20*0.89=17.8
SO ALTITUDE=17.8
AP/DA=COS(63)
AP=DA*COS(63)=20*0.4545=9.1
SO CD =AB-2*AP SINCE TRAPEZOID IS ISOCELLES..
CD=40-2*9.1=21.8
AREA=(1/2)17.8(40+21.8)=550 SQUARE UNITS.